OpenStudy (anonymous):

If sin (x+20degrees)=cos x, find the value of x in degrees?

OpenStudy (anonymous):

the value of sin and cos add up to 90

OpenStudy (anonymous):

we know sin 45 = cos 45

OpenStudy (anonymous):

so....sin(.x+20)+cosx=90?

OpenStudy (anonymous):

x+20 = 45 x = 25

OpenStudy (anonymous):

what? x=25?? how?

OpenStudy (raden):

well, just remember one of trigonometry identity is sin(90-x) = cosx so, convert cosx to sin(90-x) sin (x+20)=cos x sin (x+20)=sin(90 - x) so, x+ 20 = 90 - x 2x = 70 x = 35 degrees